\begin{thebibliography}{1}\footnotesize
\providecommand{\url}[1]{\texttt{#1}}
\providecommand{\urlprefix}{URL }

\bibitem{MR0138616}
P.~T. Bateman and R.~M. Stemmler, Waring's problem for algebraic number fields
  and primes of the form {$(p^{r}-1)/(p^{d}-1)$}, \textit{Illinois J. Math.}
  \textbf{6} (1962), 142--156.

\bibitem{bgm}
M.~A. Bennett, B.~Garbuz, and A.~Marten, Goormaghtigh's equation: small
  parameters, \textit{Publ. Math. Debrecen} \textbf{96} (2020).

\bibitem{bgk}
M.~A. Bennett, A.~Gherga, and D.~Kreso, An old and new approach to
  {G}oormaghtigh's equation, \textit{Trans. Amer. Math. Soc.} \textbf{373}
  (2020)(8), 5707--5745.

\bibitem{fgg}
M.~Filaseta, J.~Grantham, and H.~Graves, Only finitely many $s$-{C}ullen
  numbers are repunits for a fixed $s\ge 2$, preprint,
  \url{https://arxiv.org/abs/2112.04935}.

\bibitem{Goormaghtigh}
R.~Goormaghtigh, \textit{L'Interm\'ediarie de Math\'ematiciens}  (1917), 88.

\bibitem{Masser}
D.~W. Masser, Abcological anecdotes, \textit{Mathematika} \textbf{63} (2017),
  713--714.

\bibitem{Oes}
J.~Oesterl\'{e}, Nouvelles approches du ``th\'{e}or\`eme'' de {F}ermat,
  \textit{Ast\'{e}risque} \textbf{161-162} (1988), 165--186, s\'{e}minaire
  Bourbaki, exp. No. 694.

\bibitem{brazil}
B.~Schott, Les nombres br\'esiliens, \textit{Quadrature}  (2010)(76), 30--38.

\end{thebibliography}

\documentclass[10pt]{article}
\usepackage{amssymb,latexsym,amsmath,epsfig,amsthm} 
\usepackage{url}

\makeatletter

\renewcommand\section{\@startsection {section}{1}{\z@}
{-30pt \@plus -1ex \@minus -.2ex}
{2.3ex \@plus.2ex}
{\normalfont\normalsize\bfseries\boldmath}}

\renewcommand\subsection{\@startsection{subsection}{2}{\z@}
{-3.25ex\@plus -1ex \@minus -.2ex}
{1.5ex \@plus .2ex}
{\normalfont\normalsize\bfseries\boldmath}}

\renewcommand{\@seccntformat}[1]{\csname the#1\endcsname. }

\makeatother

\newtheorem{theorem}{Theorem}

\newtheorem{corollary}{Corollary}

\theoremstyle{definition}

\newtheorem{conjecture}{Conjecture}

\begin{document}
\newcommand{\rad}{\textrm{rad}}
\def\uplim{{700}}

\begin{center}
\uppercase{\bf No new Goormaghtigh primes up to $10^{\uplim}$}
\vskip 20pt
{\bf Jon Grantham}\\ 
{\small Institute for Defense Analyses/Center for Computing Sciences, Bowie, Maryland}\\
{\tt grantham@super.org}\\ 
\end{center}
\vskip 20pt


\centerline{\bf Abstract}
\noindent
The Goormaghtigh conjecture states that the only two numbers which have
two non-trivial representations as repunits are $31$ and $8191$. We call 
such a prime number a {\it Goormaghtigh prime}.
We show that there are no other Goormaghtigh primes less than $10^{\uplim}$.

\pagestyle{myheadings}
\thispagestyle{empty}
\baselineskip=12.875pt
\vskip 30pt

\section{Introduction}

Recall that a repunit is a positive integer whose digital representation in some base $b$ consists only of $1$s. Goormaghtigh \cite{Goormaghtigh} observed that the only numbers up to $10^5$ with
two non-trivial representations as repunits are $31$ (bases $2$ and $5$) and $8191$ (bases $2$ and $90$). The conjecture
that these are the only two such numbers has become known as the ``Goormaghtigh
Conjecture.'' More precisely, 
\begin{conjecture}
The only solutions to 
\begin{equation}
\label{eq:goorm}
N=\frac{x^m-1}{x-1}=\frac{y^n-1}{y-1}
\end{equation} with
integers $y>x\ge 2$, $n>m>2$ are $N=31$ and $N=8191$.
\end{conjecture}

We exclude the case $m=2$ in Conjecture 1 because every integer $N$ is a length-$2$ repunit in base $N-1$.
It is currently unknown whether the equation has finitely many solutions. Some results are known for small exponents and the case $\gcd(m-1,n-1)>1$, which we will use below.
Note that $31$ and $8191$ are both primes. We propose the terms {\it Goormaghtigh
primes} for such repunits $N$ that are primes and {\it Goormaghtigh numbers}
for any $N$ with two representations.
Prime repunits have been also studied as Brazilian primes; see \cite{brazil}. A Goormaghtigh prime can also be characterized as a Brazilian prime to two different
bases.
Bateman and Stemmler \cite{MR0138616}, using computations of Horn, noted that the only $N$ less than $1.275\times 10^{10}$ satisfying \eqref{eq:goorm} with $x$, $y$ and $N$ all prime is $31$.
In this paper, we look at a condition weaker than in the Goormaghtigh conjecture,
but stronger than in the Bateman-Stemmler question, by searching for Goormaghtigh
primes.
We use recent results of Bennett, Garbuz, and Martens \cite{bgm} as a key ingredient to reduce the computation.

\section{There Are Only Two Goormaghtigh Primes Less Than $10^{\uplim}$}

Our approach is to use lower bounds on $m$, $n$, and $y$ in
\eqref{eq:goorm} to significantly limit the number of cases less than
$10^\uplim$ that we need to check.

The following is part of Theorem 4 from \cite{bgm}.
\begin{theorem}
\label{thm:bgm}
The only solutions to Equation \eqref{eq:goorm} with $\gcd(m-1,n-1)>1$ and $m\le 50$ have $N=31$ or $N=8191$.
\end{theorem}

For Goormaghtigh primes, we must have $m$ and $n$ odd primes, and thus $m\ge 53$. 
Theorem 2 from Bennett, Gherga, and Kreso \cite{bgk} rules out further examples with $n=3$ or $n=5$ when $\gcd(m-1,n-1)>1$, which is true in the prime case. 
Theorem 3 from \cite{bgm} rules out further examples with $y\le 10^5$. 
In order to show that there are no Goormaghtigh primes below $10^{\uplim}$, we
employ the following algorithm. Let $f_k(x)=\frac{x^k-1}{x-1}$.
\begin{enumerate}
\item We perform a precomputation similar to the one in \cite{fgg}.
For a list of small primes $\{p_i\}$, we compute the values of $f_q(b)$ for all $2\le b < p_i$ and $1\le q < p_i$.
\item Generate a list of possible values of $m$. We know $m$ must be prime. From Theorem \ref{thm:bgm}, we have $m\ge 53$, and since $x>2$ in the range considered (no other known Mersenne primes are Goormaghtigh primes), $m< \log{10^{\uplim}}/\log{3}\approx 1467$.
\item For each $m$ in the above list, we examine all $x$ with $x>2$
and $x<10^{\uplim/m}$. If $x \not\equiv 1 \pmod {p_i}$, then
$f_m(x)\equiv f_{m'}(x)\pmod {p_i}$ for $m\equiv m'\pmod{p_i-1}$. We retrieve this value from the precomputation. Call this value $a_i$.
\item We examine each prime $n$ such that $m>n\ge 7$; we exclude the
case $3(m-1)=(n-1)$ by Theorem 6 of \cite{bgm}. For $n'\equiv n\pmod
{p_i-1}$, we see if $a_i$ is a possible value of $f_{n'}(y)$. To handle
the case $y\equiv 0\pmod {p_i}$, we check if $a_i\equiv 1\pmod {p_i}$.
Finally, if $y\equiv 1$, we have $f_n(y)\equiv n$, so we check if $a_i\equiv n$. If none of these conditions hold, we have shown there is no solution mod $p_i$, and thus over the integers. If there is a solution, we repeat with a different $p_i$.

\end{enumerate}

The computation up to $10^{500}$ took $129$ minutes on a single core of an Intel SP Platinum 8280 CPU running at 2.7 GHz. The computation up to $10^{\uplim}$ took approximately $480$ core-days.

Note that we have, in fact, shown a slightly stronger result, that there are no new Goormaghtigh numbers up to $10^{\uplim}$ where both exponents are prime.

\section{A Conditional Result}

Recall the abc conjecture.

\begin{conjecture}
The \textit{abc conjecture} of Oesterl\'e \cite{Oes} and Masser
\cite{Masser} states that if $a$, $b$, and $c$ are relatively prime
integers such that $a + b = c$, then for any $\epsilon >0$, only
finitely many $(a,b,c)$ fail to satisfy the inequality
\[ c < \rad(abc)^{1+\epsilon}.\]
\end{conjecture}

Carl Pomerance suggested an argument that gives the following.

\begin{theorem}
Assuming the abc conjecture, there are only finitely many Goormaghtigh numbers where neither representation is of length three or four.
\end{theorem}

\begin{proof}
We see that $y^{n+1}>x^m$, so we will use that $y^{(n+1)/m}>x$.
Rewrite \eqref{eq:goorm} as
$$(x^m-1)(y-1)=(y^n-1)(x-1).$$
Rearranging,
$$x^m(y-1)+(x-y)=y^n(x-1).$$
Let $$g=\gcd(x^m(y-1),x-y,y^n(x-1)).$$
Let $a=x^m(y-1)/g$, $b=(x-y)/g$ and $c=y^n(x-1)/g$.
Then we have $a+b=c$, and
$$\rad(abc)\le x(y-1)(y-x)y(x-1)/g<y^{3+(n+1)/m}(x-1)/g.$$
On the other hand, we have $c=y^n(x-1)/g$.
Let $\epsilon=1/4$. Then it suffices to show that for $n>m\ge 5$,
$\rad(abc)^{1+\epsilon}/c\le 1$,
or $$\left(y^{3+(n+1)/m}(x-1)/g\right)^{5/4}/(y^n(x-1)/g)\le 1.$$
This is equivalent to $$y^{15/4+(5/4)(n+1)/m-n}((x-1)/g)^{1/4}\le 1.$$ The exponent on $y$ is no greater than $-1/2$ (which is achieved when $m=5$, $n=6$). We have by necessity that $((x-1)/g)^{1/4}\le(x-1)^{1/4}<
(x-1)^{1/2}$, so we have established the conditions for the abc conjecture. Therefore there are only finitely many Goormaghtigh numbers
with $m<5$.
\end{proof}

Combining Theorem 2 with Theorem 2 from \cite{bgk}, which eliminates the
length-3 case for primes, gives the following.

\begin{corollary}
Assuming the abc conjecture, there are only finitely many Goormaghtigh primes.
\end{corollary}

\vskip20pt\noindent{\bf Acknowledgements.} The author is grateful to Mike Bennett for helpful e-mail exchanges and providing copies of code, to Hester Graves for helpful conversations and checking the Mersenne primes, to Xander Faber and Mark Morgan for comments on an earlier draft of this paper, and to Gregory Minton for suggesting ruling out solutions locally.

\end{document}